1) A particle is constrained to travel along the path. If x = (4t^2)m, where t is in seconds, determine the magnitude of the particle’s velocity and acceleration when t= 0.5 s.
2) A particle traveling along a parabolic path y = 0.25x^2. If x= (2t^2)m, where t is in seconds,
determine the magnitude of the particle’s velocity and acceleration when t = 2s.
1
Expert's answer
2020-08-27T10:40:48-0400

(i)

Given"\ x = 4t^2"

as we know that velocity is,V = "\frac{dx}{dt}"

so "\frac{dx}{dt} = 8t"

velocity when t = 0.5 ; V = "8\times0.5 = 4 m/s"

Acceleration is "\frac{d^2x}{dt^2}" which is equal to 8. Hence acceleration is 8 m/s2


(ii) Given "y = 0.25x^2 \ and \ x= 2t^2"

substituting value of x in y we get, "y = 0.25 \times[2t^2]^2"

"y= t^4"

Velocity (V) is "\frac{dy}{dt}" and acceleration (a) is "\frac{d^2y}{dt^2}"

"and \ \frac{dy}{dt} = 4t^3"

so velocity at t= 2sec; V = "4\times 2^3 = 32m/s"

and "\frac {d^2y}{dt^2}= 12t^2"

so acceleration at t= 2 sec ;

a = "12\times 2^2 = 48m/s^2"